OpenStudy (anonymous):

What is the direct variation equation if y varies directly with x and y = –9 when x = –45? A: y = 1/5x B: y = -5x C: y = 5x D: y = -1/5x

OpenStudy (anonymous):

y = kx Plug in what we know: 'y = -9' 'x = -45' -9 = k(-45) -9 = -45k Divide -45 to both sides.

OpenStudy (anonymous):

Wait, so -9 divided by -45?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

No..idk how you got that..try again. -9 / -45

OpenStudy (anonymous):

my calculator glitched, sorry

OpenStudy (anonymous):

It's fine

OpenStudy (anonymous):

0.2?

OpenStudy (anonymous):

Yes, and 0.2 is the same thing as 2/10 which simplifies to 1/5. So the constant of variation is 1/5. So we write it like: y = 1/5x

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

No problem.